Test the series ∞ Σ n=1 (-1)^n-1. 3/(5n-2) for absolute and conditional convergence

1
Expert's answer
2022-02-22T11:19:06-0500

ANSWER The series "\sum_{n=1}^{\infty}\frac{{3\cdot \left (-1 \right )}^{n-1}}{5n-2}" is conditionally convergent .

EXPLANATION.

First we check absolute convergence . "\sum_{n=1}^{\infty}\ \left | \frac{{3\cdot \left (-1 \right )}^{n-1}}{5n-2} \right |=\sum_{n=1}^{\infty}\frac {3 }{5n-2}" . Let "a _{n}=\frac{3}{5n-2}."

"\lim_{n \rightarrow\infty}\frac{\frac{3}{5n-2}}{\frac{1}{n}}=\lim_{n \rightarrow\infty}\frac{3n}{5n-2}=\lim_{n \rightarrow\infty}\frac{3 }{5 -\frac{2}{n}}=\frac{3}{5}" . Since "\lim_{n \rightarrow\infty}\frac {a_{n}} {\frac{1}{n}}=\frac{3}{5}>0" and the series "\sum_{n=1}^{\infty}\frac{1}{n}" diverges ,then by the Limit Comparison Test "\sum_{n=1}^{\infty}\frac {3 }{5n-2}" diverges. Therefore , the series does not converge absolutely.

The series "\sum_{n=1}^{\infty}\frac{{3\cdot \left (-1 \right )}^{n-1}}{5n-2}" is alternating. Check the two conditions: "1) \lim_{n\rightarrow\infty}a_{n}=\lim_{n\rightarrow\infty}\frac{3} {5n-2}=0" . "2)a _{n+1} \leq a_{n}" , because "5\cdot\left ( n+1 \right ) -2 \ge 5 n-2\\" so

"\frac{3}{5\cdot\left ( n+1 \right ) -2}\le \frac{3}{5n-2}" .

Since the two conditions of the Alternating Series Test are satisfied , "\sum_{n=1}^{\infty}\frac{{3\cdot \left (-1 \right )}^{n-1}}{5n-2}" is conditionally convergent by the Alternating Series Test .